A glass container is knocked off a shelf at a height of 1.4 meters above the floor beside you. It takes you 0.35 seconds to react to the falling object and reach out your hand in an attempt to catch the container. Will you be able to reach the container before it hits the floor?
1
Expert's answer
2019-04-12T09:39:19-0400

The time of flight is


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(1.4)}{9.8}}=0.29 s."

It is less than 0.35 seconds. Thus, we will not be able to reach the container before it hits the floor.